Please help!!
(x+8)^2 +15(x+8) +54 = 0

Answer by lwsshak3(11628)

(x+8)^2 +15(x+8) +54 = 0
let u=(x+8)
u^2+15u+54=0
(u+9)(u+6)=0
..
u=-9=x+8
x=-17
..
u=-6=x+8
x=-14
..
Check:
f(-17)=(x+8)^2 +15(x+8) +54
=(-17+8)^2 +15(-17+8) +54=0
..
f(-14)=(x+8)^2 +15(x+8) +54
=(-14+8)^2+15(-14+8)+54=0
check ok